# The inherent goodness of well educated intelligence

### v6

@@ Appendix A Simple Beauty and the Crazy Ones
The correct theory is the “simply beautiful” theory as postulated by Physics Nobel Laureates Murray Gell-Mann and Paul Dirac. Simple can be Frank Lloyd Wright’s Fallingwater, Phillip Johnson’s postmodern Pennzoil building, Steve Jobs’ electronic devices, a Danish Bang & Olufsen stereo, a Piet Mondrian painting, or self organized fluid flow constrained by topological invariants (see Fig. 9). There is also the simple beauty of textures, that is Benoit Mandelbrot’s fractals, a Jackson Pollock painting, a Brazilian favela, the chaos boards of Carrie Mathison in “Homeland” or John Nash in “A Beautiful Mind”, or Andrey Kolmogorov’s well developed turbulence (see Fig. 10). The later is as simple as the former. There is a high degree of organization or correlation in the later. Our visual and audio cortex, along with the Convolutional Neural Networks (CNNs) of modern AI are very good at identifying the simple beauty or symmetry in these complex images or signals.
@@
Figure 11: Jackson Pollock action painting. (a) An action painting by Jackson Pollock, Autumn Rhythm (Number 30). (b) Jackson Pollock action painting Autumn Rhythm, photograph by Rudolph Burkhardt (Video in Italian), and a film by Hans Namuth of Jackson Pollock painting Autumn Rhythm (“Autumn Rhythm in Summer”).
While it is easy to recreate a minimalistic painting like a Mondrian in the collective or field domain, a maximalistic painting like a Pollock is much easier to recreate in the process or dynamical domain of the puppet master or painter. The deep deconvolution of the HST and the MLP decoder should be used to determine the motions or symmetries of the artist’s movements, then the movements of the artist should be simulated and deeply convolved into the field domain of the painting. While it is not necessary to do this process based recreation for a minimalistic painting, it can be used to generate a painting in the style of that artist that was not painted by that artist, but easily could have been painted by that artist. This is a hallucination, not an exact copy. This is what OpenAI’s DALL-E has done in Fig. 8 to create the illustrations for the methods of control, and, in an homage to the action painting of Pollock, a series of “rhythms” for each of the seasons, shown in Fig. 12.

### v8

@@ Appendix A Simple Beauty and the Crazy Ones
The correct theory is the “simply beautiful” theory as postulated by Physics Nobel Laureates Murray Gell-Mann (Johnson, 2000) and Paul Dirac (Salam, 1990). Simple can be Frank Lloyd Wright’s Fallingwater, Phillip Johnson’s postmodern Pennzoil building, Steve Jobs’ electronic devices, a Danish Bang & Olufsen stereo, a Piet Mondrian painting, or self organized fluid flow constrained by topological invariants (see Fig. 10). There is also the simple beauty of textures, that is Benoit Mandelbrot’s fractals, a Jackson Pollock painting, a Brazilian favela, the chaos boards of Carrie Mathison in “Homeland” or John Nash in “A Beautiful Mind”, or Andrey Kolmogorov’s well developed turbulence (see Fig. 11). The later is as simple as the former. There is a high degree of organization or correlation in the later. Our visual and audio cortex, along with the Convolutional Neural Networks (CNNs) of modern AI are very good at identifying the simple beauty or symmetry in these complex images or signals.
@@
Figure 12: Jackson Pollock action painting. (a) An action painting by Jackson Pollock, Autumn Rhythm (Number 30). (b) Jackson Pollock action painting Autumn Rhythm, photograph by Rudolph Burkhardt (YouTube video in Italian), and a film by Hans Namuth of Jackson Pollock painting Autumn Rhythm (Vimeo video of “Autumn Rhythm in Summer”).
While it is easy to recreate a minimalistic painting like a Mondrian in the collective or field domain, a maximalistic painting like a Pollock is much easier to recreate in the process or dynamical domain of the puppet master or painter. The deep deconvolution of the HST and the MLP decoder should be used to determine the motions or symmetries of the artist’s movements, then the movements of the artist should be simulated and deeply convolved into the field domain of the painting. While it is not necessary to do this process based recreation for a minimalistic painting, it can be used to generate a painting in the style of that artist that was not painted by that artist, but easily could have been painted by that artist. This is a hallucination, not an exact copy. This is what OpenAI’s DALL-E has done in Fig. 13, creating the illustrations in an homage to the action painting of Pollock, a series of “rhythms” for each of the seasons.